OpenStudy (anonymous):

(x+2)^2-(x+7)^2 Using a^2-b^2=(a+b)(a-b) Answer from the book is -5(2x+9) I cannot get there from here. If anyone has time, could you please explain how to arrive at that answer?

OpenStudy (zzr0ck3r):

\((x+2)^2-(x+7)^2=\\x^2+4x+4-(x^2+14x+49=\\x^2+4x+4-x^2-14x-49=\\-10x-45=-5(2x+9)\)

OpenStudy (anonymous):

\[\left( x+2 \right)^2-\left( x+7 \right)^2=\left\{ x+2+x+7\right\}\left\{ x+2-\left( x+7 \right) \right\}\] \[=(2x+9)(x+2-x-7)=-5(2x+9)\]

OpenStudy (zzr0ck3r):

weird use of set brackets:)

OpenStudy (anonymous):

Thank you both! Why can't I give best response to both? What a crock, you both helped.

OpenStudy (zzr0ck3r):

give it to @surjithayer , his\hers was pretty

OpenStudy (anonymous):

Done! Thank you both again, its whipping me but with help Im getting the hang of it.

OpenStudy (anonymous):

yw. i am a man.

OpenStudy (zzr0ck3r):

I should delete my comment, then it would make that last statement seem strange.